英語 (English)

基盤英語 (Basic English) (医 ((医 C)1 年))

座喜純 · 准教授/ 大学院ソシオ・アーツ・アンド・サイエンス研究部 1 単位前期月 7 ・ 8

(平成 19 年度以前の授業科目: 『基盤英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (1)』 )

【授業の目的】 This classs designed to give a good assistance in developing Basic English abilities in general. Especially, It will focus on Grammar, Listening, Reading and Writing skills.

【授業の概要】 Classes will be running in ENGLISH language only to give more practice chance to everyone. There is no specific text book but there will be some other media will be selected up to date. Listening and Reading practices will take place during the class hour. Grammar guidance, Listening and Reading practices will take place during the class hour.

【キーワード】 [キーワード] 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【到達目標】 You will gain more skills as much as you do efforts to practice. 【授業の計画】

1. Overall Guidance

2. Clip 1

3. Clip 2

4. Clip 3

5. Clip 4

6. Clip 5

7. Clip 6

8. Clip 7

9. Clip 8

10. Clip 9

11. Clip 10

12. Clip 11

13. Clip 12

14. Clip 13

15. Clip 14

16. Review 【教科書】 No specific text book but there will be some other media will be selected

up to date. 【参考書等】 [参考資料] 【成績評価の方法】 Attendance and Participation in class 【再試験の有無】 No make-up examination for this course. 【受講者へのメッセージ】 Asking questions in class is very important.

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220824 【連絡先 (オフィスアワー・研究室・ E メールアドレス)】 ⇒ 座喜 .

English

Basic English (医 ((医 C)1 年))

Jun Zaki · Associate Professor / Institute of Socio-Arts and Sciences 1 unit 前期月 7 ・ 8

(平成 19 年度以前の授業科目: 『基盤英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (1)』 )

Target⟩ This classs designed to give a good assistance in developing Basic English abilities in general. Especially, It will focus on Grammar, Listening, Reading and Writing skills.

Outline⟩ Classes will be running in ENGLISH language only to give more practice chance to everyone. There is no specific text book but there will be some other media will be selected up to date. Listening and Reading practices will take place during the class hour. Grammar guidance, Listening and Reading practices will take place during the class hour.

�������⟩ [キーワード] Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Goal⟩ You will gain more skills as much as you do efforts to practice.

Schedule⟩

1. Overall Guidance

2. Clip 1

3. Clip 2

4. Clip 3

5. Clip 4

6. Clip 5

7. Clip 6

8. Clip 7

9. Clip 8

10. Clip 9

11. Clip 10

12. Clip 11

13. Clip 12

14. Clip 13

15. Clip 14

16. Review Textbook⟩ No specific text book but there will be some other media will be

selected up to date. Reference⟩ [参考資料] Evaluation Criteria⟩ Attendance and Participation in class Re-evaluation⟩ No make-up examination for this course. Message⟩ Asking questions in class is very important.

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220824 Contact (Office-Hour, Room, E-mail)⟩

⇒ Zaki .